OpenStudy (anonymous):

If the pyramid in the previous problem had a Volume of 125,000ft3, length of 75 fee, and height of 100 feet, how wide was it?

OpenStudy (callisto):

@NeedMuchHelp What is the formula you've got for your previous question?

OpenStudy (anonymous):

V = 1/3 lwh

OpenStudy (callisto):

What about the answer to your previous question, the one with w (width) as the subject?

Parth (parthkohli):

Hmm. \( \color{Black}{\Rightarrow 125,000 = {1 \over 3} \times 75 \times 100 \times w}\) Can you solve this equation?

OpenStudy (anonymous):

Put the values of V, l and h in the equations. 125000= 1/3 * 75 * 100 * w Now can you solve it?

OpenStudy (anonymous):

im confused on how to.

OpenStudy (callisto):

@NeedMuchHelp May I know what you've got in your previous question? The answer to your previous question helps a lot in doing this question.

OpenStudy (anonymous):

first divide by 100 on both sides of the equation. It becomes 1250 = 1/3 * 75 *w 1/3*75 = 25 So, 1250 = 25 * w Now do you know how to do it?

OpenStudy (anonymous):

yes